# Organizing Unstructured Image Collections using Natural Language

### v4

@@ Appendix D Further Details of Evaluation Protocol
Appendix E Further Implementation and Prompt Details
In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in App. E.1 and the Semantic Grouper in App. E.2.
E.1 Further Details of Criteria Proposer
@@
In Tab. 7, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B (Li et al., 2025). Given a target image set, we first randomly shuffle the images and divide them into disjoint subsets, each containing 64 images. Each subset is then stitched into an $8\times 8$ image grid, treated as a single image, and fed into the MLLM. For each subset, the MLLM is prompted to propose 5 distinct grouping criteria for organizing the images within that subset, using the prompt shown in Tab. 7. After iterating through all subsets, we take the union of the criteria proposed for each subset as the discovered criteria for the target image set. Finally, we deduplicate the discovered criteria and accumulate them into the criteria pool.
@@
In Tab. 8, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B (Meta, 2024a). For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 (Radford et al., 2021), to assign 10 related natural language tags to each image. These tags are selected from the WordNet (Miller, 1995) vocabulary, which contains 118k English synsets, and represent the semantic content of the images. We employ the standard prompt “A photo of {concept}” provided by CLIP for image tagging. Next, we embed the assigned tags into the prompt shown in Tab. 8 to carry the semantics of the entire image set and query the LLM to propose grouping criteria. The criteria proposed by the LLM are then added to the criteria pool. Note that in this case, we embed the tags for the entire dataset into a single prompt for criteria proposal, without reaching the LLM context length limits (e.g., 128k for Llama-3.1-8B) for the datasets used in our experiments. However, for larger datasets, it may be necessary to split the dataset into subsets, prompt the LLM for each subset, and use the union of the proposed criteria as the final output.
@@
We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B (Liu et al., 2024b) in Tab. 9, and the prompt for the LLM Llama-3.1-8B (Meta, 2024a) in Tab. 10. Specifically, we first use the MLLM with a general prompt to generate detailed descriptions for each image in the target dataset, effectively translating the visual information into natural language. The generated captions are then randomly shuffled and split into disjoint subsets, each containing 400 captions. Next, we embed the captions from each subset into the prompt shown in Tab. 10 and use it to query the LLM to propose grouping criteria for the images represented by the captions. After iterating through all subsets, we take the union of the proposed criteria across subsets as the discovered criteria for the target image set. Finally, we deduplicate these criteria and add them to the criteria pool. Due to the context window limitations of LLMs, embedding all captions into a single prompt is infeasible. To address this, we limit each subset to 400 captions, which results in approximately 115k tokens per subset. This strategy allows us to remain within the context length limits of modern LLMs (e.g., 128k tokens for both Llama-3.1 and GPT-4o) while maximizing the number of samples per query to effectively propose clustering criteria.
@@
Since the Proposers operate on subsets, the criteria accumulated in $\mathcal{R}$ may include redundant or noisy entries. Therefore, we further include a criteria poop refinement process. Specifically, we input all initially proposed criteria from $\mathcal{R}$ into a LLM, prompting it to consolidate similar criteria and discard noisy ones. This process refines and updates the criteria pool $\mathcal{R}$ for next stage.
In Tab. 11, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B (Meta, 2024a). Since the accumulated criteria pool $\widetilde{\mathcal{R}}$ may contain highly similar or noisy clustering criteria, we embed the criteria from the pool into the prompt shown in Tab. 11 and ask the LLM to merge similar criteria and rephrase their names to enhance clarity. This process yields a refined set of grouping criteria, which is then passed to the next stage for image grouping.
E.2 Further Details of Semantic Grouper
@@
Task Instruction Therefore, I need your assistance in designing a prompt for the Visual Question Answering (VQA) model to help it identify the "{CRITERION}" category in a given image at three different granularity. Please help me design and generate this prompt using the following template: "Question: [Generated VQA Prompt Question] Answer (reply with an abstract, a common, and a specific category name, respectively):". The generated prompt should be simple and straightforward.
@@
In the tag-based grouper, we begin by embedding the target criterion into the prompt from Tab. 13 to generate criterion-specific tags at a middle granularity. To enhance the diversity and coverage of the tags, we query the LLM 10 times and take the union of the generated tags after deduplication as candidates. Following the SHiNe framework (Liu et al., 2024d), for each middle-grained tag, we further embed it into the prompts from Tab. 14 and Tab. 15 to generate 3 super-categories (coarse-grained) and 10 sub-categories (fine-grained) for each tag.
After generating coarse and fine-grained categories for all middle-grained tags, we take the union of the super-categories as the coarse-grained tag candidates and the union of the sub-categories as the fine-grained tag candidates. Lastly, we use the open-vocabulary tagger CLIP ViT-L/14 to assign the most relevant tags to each image based on cosine similarity, using candidates from each granularity level. After tagging all the images, we group those sharing the same tag into clusters, yielding the clustering result. Note that we do not utilize lexical databases such as WordNet (Miller, 1995) or ConceptNet (Speer et al., 2017) for tag generation, as they do not support free-form input and may not capture certain discovered criteria.
@@
Output Instruction Please respond in the following format within five words: "*Answer*". Do not talk about the description and do not respond long sentences. The answer should be within five words.
@@
Sub-task Instruction Follow these steps to generate the hierarchy: Step 1 - Understand the provided initial list of "{CRITERION}" categories. The following three-level class hierarchy generation steps are all based on the provided initial list. Step 2 - Generate a list of abstract or general "{CRITERION}" categories as the first level of the class hierarchy, covering all the concepts present in the initial list. Step 3 - Generate a list of middle-grained "{CRITERION}" categories as the second level of the class hierarchy, in which the middle-grained categories are the subcategories of the categories in the first level. The categories in the second-level are more specific than the first level but should still cover and reflect all the concepts present in the initial list. Step 4 - Generate a list of more specific fine-grained "{CRITERION}" categories as the third level of the class hierarchy, in which the categories should reflect more specific "{CRITERION}" concepts that you can infer from the initial list. The categories in the third-level are subcategories of the second-level. Step 5 - Output the generated three-level class hierarchy as a JSON object where the keys are the level numbers and the values are a flat list of generated categories at each level, structured like: { “level 1": [“categories”], “level 2": [“categories”], “level 3": [“categories”] }
@@
We first present the MLLM prompt used for LLaVA-NeXT-7B (Liu et al., 2024b) to generate criterion-specific captions in Tab. 16. Following this, we present the LLM Llama-3.1-8B prompts used in the caption-based grouper for the Initial Naming step in Tab. 17, the Multi-granularity Cluster Refinement step in Tab. 18, and the Final Assignment step in Tab. 19.
@@
Appendix F Further Details of the Compared Methods
@@
Implementation details of MMaP and MSub (Yao et al., 2024, 2025): We closely followed the training configuration outlined in the original MMaP and MSub paper. Specifically, GPT-turbo-2024-04-09 was used as the LLM to generate reference words for each dataset. We then prompt-tuned CLIP-ViT/B32 using Adam with a momentum of 0.9, training the model for 1,000 epochs for each criterion across all datasets. Hyperparameters were optimized according to the loss score of MMaP, with the learning rate searched in {0.1, 0.05, 0.01, 0.005, 0.001, 0.0005}, weight decay in {0.0005, 0.0001, 0.00005, 0.00001, 0}, $\alpha$ and $\beta$ in {0.0, 0.1, 0.2, …, 1.0}, and $\lambda$ fixed at 1 for all experiments. After training, KMeans, with the ground-truth number of clusters, was applied for each criterion and dataset to perform clustering.
Appendix G Further Quantitative Experimental Results
In this section, we present additional numerical experiment results to supplement the figures in the main paper. In § G.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework. In § G.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks. Furthermore, we include expanded results comparing our framework to prior criteria-conditioned clustering methods. Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in § G.4.
G.1 Further Results for Criteria Proposer Study
@@
Although captions generated by the MLLM may exhibit some information loss (e.g., ignoring small objects or attributes) (He et al., 2024) and hallucinations (e.g., introducing objects not present in the images) Liu et al. (2024a), these issues generally occur at the object or fine-grained attribute level. However, when reasoning about grouping criteria for OpenSMC task, the focus is on identifying general thematic elements shared across the image set. As a result, these minor inconsistencies in the captions do not hinder the LLM in our framework from effectively reasoning about grouping criteria, helping the Caption-based Proposer to achieve the best performance among all the studied design choices.
@@
G.2 Further Results for Semantic Grouper Study
@@
G.3 Further Comparative Results with TCMC Methods
@@
G.4 Further Results for Multi-granularity Clustering Study.
@@
Appendix H Qualitative Analysis
@@
Appendix I Failure Case Analysis
@@
Finally, we observed that our method sometimes fails to distinguish subtle, fine-grained differences between images, leading to incorrect labels. For example, as shown in Figure 16, “Edamame” or “Pho” are typical dishes from China, Vietnam, and Japan, but they may be presented differently depending on the cuisine. The “Edamame” shown in Figure 16 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine. This oversight of fine-grained details could be improved by employing a more advanced prompting strategy (Liu et al., 2024e).
@@
Appendix J Further Study on Multi-granularity Clustering
@@
Appendix K Study on Handling Invalid Criteria
@@
To this end, we design and conduct a control experiment using the Fruit-2c dataset (Muresan and Oltean, 2018), where we artificially introduced two “hallucinated” invalid grouping criteria (False Positives), Action and Clothing Style, into the refined criteria pool. These invalid criteria were then used in the subsequent grouping process to evaluate their effect on our system. We apply the main Caption-based Grouper to group fruit images based on these “hallucinated” criteria.
@@
Appendix L Study on Model Hallucination and Bias
@@
Model hallucination. LLM hallucination (Wang et al., 2024c) typically occurs when LLMs are tasked with complex queries requiring world knowledge or factual information—for instance, answering a question like "Who was the 70th president of the United States?" might lead to a fabricated response. However, in our system, the use of LLMs is fully grounded in the visual descriptions (tags or captions) of the images. Consequently, the LLM output is strongly constrained to analyzing these visual descriptions, significantly reducing the likelihood of hallucination. That said, LLM hallucination can still have mild effects on clustering results. For example, as discussed in the failure case analysis in Appendix I, the LLM incorrectly grouped “Korean bibimbap” and “Vietnamese rice noodles” under “Chinese cuisine” (see Figure 16). MLLMs also play a crucial role in our system, as they are responsible for translating images into text for subsequent processing steps. MLLM hallucination (Wang et al., 2024c) typically involves incorrectly identifying the existence of objects, attributes, or spatial relationships within an image. However, since our proposed system operates at the dataset level rather than on a per-image basis, it is largely insensitive to such hallucinations, especially at the fine-grained visual detail level. Moreover, as our system is training-free, it can be further enhanced with LLM or MLLM hallucination mitigation techniques, such as the Visual Fact Checker (Ge et al., 2024), which we leave as a direction for future work.
@@
Model Bias. Foundation models such as LLMs and MLLMs are known to inherit biases from their training data (Bommasani et al., 2021). In our system, we addressed potential biases using Hard Positive Prompting techniques: i) MLLM Bias Mitigation: The MLLM is further prompted to generate criterion-specific captions that focus solely on describing the criterion-related content in each image. This approach constrains the MLLM from generating irrelevant content influenced by inherent biases; ii) LLM Bias Mitigation: Similarly, when prompting the LLM to assign image captions to clusters, we condition it to concentrate exclusively on the Criterion depicted in each image (see Tab. 17).
@@
To validate the effectiveness of these bias mitigation techniques, we conducted a fair clustering experiment. Specifically, following Kwon et al. (2024), we sampled images for four occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET (Gustafson et al., 2023) dataset, which contains images from 52 occupations. For each occupation, we selected 10 images of men and 10 images of women, totaling 80 images, ensuring a ground-truth gender proportion disparity of 0% for each occupation. Using our main $\mathcal{X}$-Cluster system, we grouped these images based on the criterion Occupation using three bias mitigation strategies: i) No mitigation: using general descriptions from the MLLM for LLM grouping; ii) Our default hard positive prompting strategy: using criterion-specific captions from the MLLM for LLM grouping; and iii) Our default strategy with additional negative prompt: adding a simple negative prompt, “Do not consider gender,” to both the MLLM captioning and LLM grouping prompts.
@@
Appendix M Computational Cost Analysis
The proposed $\mathcal{X}$-Cluster framework is training-free, requiring only inference processes. Specifically, our main framework (Caption-based) requires up to 31 GB of GPU memory to operate. All experiments reported in the paper were conducted on 4 Nvidia A100 40GB GPUs. In Tab. 33, we provide a detailed analysis of the computational efficiency of our main $\mathcal{X}$-Cluster framework (Caption-based Proposer and Caption-based Grouper) on the COCO-4c benchmark (5,000 images with four criteria) across various hardware configurations. For these experiments, we used LLaVA-NeXT-7B (Liu et al., 2024b) as the MLLM and Llama-3.1-8B (Meta, 2024a) as the LLM.
As shown in Tab. 33, organizing 5,000 images based on all four discovered criteria can be completed by $\mathcal{X}$-Cluster in 29.1 hours on a single A100 GPU or 16.7 hours on a single H100 GPU. More importantly, most steps in our framework, such as per-image captioning and per-caption cluster assignment, are parallelizable across multiple GPUs, significantly accelerating the process. Therefore, when parallelizing the framework on 4 A100 or H100 GPUs, we achieve approximately a 4$\times$ speedup, reducing computational time to 7.6 hours on 4 A100 GPUs and 4.3 hours on 4 H100 GPUs.
@@
Appendix N System Sensitivity Analysis of Various MLLMs and LLMs
@@
Specifically, in Figure 18(a), we first fix the LLM in our system to Llama-3.1-8B (Meta, 2024a) and assess the influence of various MLLMs: GPT-4V (OpenAI, 2023), BLIP-3-4B (Xue et al., 2024), and LLaVA-NeXT-7B (Liu et al., 2024b). Next, in Figure 18(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo (OpenAI, 2023), GPT-4o (OpenAI, 2024), Llama-3-8B (Meta, 2024b), and Llama-3.1-8B.
@@
As for the LLMs, as depicted in Figure 18(b), despite GPT-4-turbo showing marginally superior performance, the open-source Llama-3.1-8B achieves similar results across benchmarks, making it our default LLM. Notably, except for the Card-2c dataset, system performance remains largely consistent regardless of the power of the LLM. This consistency suggests that the reasoning task for OpenSMC, given the capabilities of modern LLMs to tackle complex problems (Street et al., 2024), is relatively straightforward.
Appendix O Study on Fine-grained Image Collections
Image collections may include fine-grained grouping criteria, such as Bird species in bird photography. Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations (Zhang et al., 2014; Vedaldi et al., 2014; He and Peng, 2017). This requires the model to detect subtle visual distinctions to accurately infer cluster names and guide the grouping process. The straightforward captioning process in our current framework may not fully capture these subtle visual nuances. However, the modular design of our framework allows for seamless integration of advanced cross-modal chain-of-thought (CoT) prompting strategies to address this issue.
We demonstrate this by enhancing our Caption-based Grouper with FineR (Liu et al., 2024e), a cross-modal CoT prompt method specifically designed for fine-grained visual recognition. When the proposer identifies fine-grained criteria, such as Bird species, the framework switches to a FineR-enhanced captioning strategy that provides more detailed attribute descriptions, such as “Wing color: Blue-grey,” to enrich the captions and capture per-attribute visual characteristics to better support the subsequent substructure uncovering process.
Table 34: Study of substructure discovery for fine-grained criteria. We report clustering accuracy (CAcc) and semantic accuracy (SAcc) as percentages (%). The pseudo upper-bound (UB) performance is obtained using CLIP (Radford et al., 2021) ViT-L/14 in a zero-shot transfer setting with the ground-truth class names. †: We compare with FineR (Liu et al., 2024e) without its post-class name refinement step to ensure a fair comparison.

### v5

@@ Appendix E Further Details of Evaluation Protocol
Appendix F Further Implementation and Prompt Details
In this section, we provide detailed descriptions of the exact prompts used in our framework, along with additional implementation details for the proposed Criteria Proposer in Supp. F.1 and the Semantic Grouper in Supp. F.2.
Figure 1: All three variants of the proposed $\bm{\mathcal{X}}$-Cluster framework. We explore different design choices for both the Criteria Proposer (left) and the Semantic Grouper (right), and designate the best-performing Caption-based system (marked with Refer to caption) as the main $\mathcal{X}$-Cluster configuration in our experiments.
Further, we present a system overview illustrating all three variants of our Criteria Proposer and Semantic Grouper, namely the caption based (main), tag based, and image based versions, in Figure 1.
F.1 Further Details of Criteria Proposer
@@
In Tab. 5, we present the exact prompt used in the image-based proposer for querying the MLLM LLaVA-NeXT-Interleave-7B [li2024llava]. Given a target image set, we first randomly shuffle the images and divide them into disjoint subsets, each containing 64 images. Each subset is then stitched into an $8\times 8$ image grid, treated as a single image, and fed into the MLLM. For each subset, the MLLM is prompted to propose 5 distinct grouping criteria for organizing the images within that subset, using the prompt shown in Tab. 5. After iterating through all subsets, we take the union of the criteria proposed for each subset as the discovered criteria for the target image set. Finally, we deduplicate the discovered criteria and accumulate them into the criteria pool.
@@
In Tab. 6, we present the exact prompt used in the tag-based proposer for querying the LLM Llama-3.1-8B [meta2024llama31]. For a given target image set, we first utilize an open-vocabulary tagger, CLIP ViT-L/14 [radford2021learning], to assign 10 related natural language tags to each image. These tags are selected from the WordNet [miller1995wordnet] vocabulary, which contains 118k English synsets, and represent the semantic content of the images. We employ the standard prompt “A photo of {concept}” provided by CLIP for image tagging. Next, we embed the assigned tags into the prompt shown in Tab. 6 to carry the semantics of the entire image set and query the LLM to propose grouping criteria. The criteria proposed by the LLM are then added to the criteria pool. Note that in this case, we embed the tags for the entire dataset into a single prompt for criteria proposal, without reaching the LLM context length limits (e.g., 128k for Llama-3.1-8B) for the datasets used in our experiments. However, for larger datasets, it may be necessary to split the dataset into subsets, prompt the LLM for each subset, and use the union of the proposed criteria as the final output.
@@
We present the prompt used in the caption-based proposer for the MLLM LLaVA-NeXT-7B [liu2024llava] in Tab. 7, and the prompt for the LLM Llama-3.1-8B [meta2024llama31] in Tab. 8. Specifically, we first use the MLLM with a general prompt to generate detailed descriptions for each image in the target dataset, effectively translating the visual information into natural language. The generated captions are then randomly shuffled and split into disjoint subsets, each containing 400 captions. Next, we embed the captions from each subset into the prompt shown in Tab. 8 and use it to query the LLM to propose grouping criteria for the images represented by the captions. After iterating through all subsets, we take the union of the proposed criteria across subsets as the discovered criteria for the target image set. Finally, we deduplicate these criteria and add them to the criteria pool. Due to the context window limitations of LLMs, embedding all captions into a single prompt is infeasible. To address this, we limit each subset to 400 captions, which results in approximately 115k tokens per subset. This strategy allows us to remain within the context length limits of modern LLMs (e.g., 128k tokens for both Llama-3.1 and GPT-4o) while maximizing the number of samples per query to effectively propose clustering criteria.
@@
In Tab. 9, we present the exact prompt used for criteria pool refinement when querying the LLM Llama-3.1-8B [meta2024llama31]. Since the accumulated criteria pool $\widetilde{\mathcal{R}}$ may contain highly similar or noisy clustering criteria, we embed the criteria from the pool into the prompt shown in Tab. 9 and ask the LLM to merge similar criteria and rephrase their names to enhance clarity. This process yields a refined set of grouping criteria, which is then passed to the next stage for image grouping.
F.2 Further Details of Semantic Grouper
@@
Task Instruction Therefore, I need your assistance in designing a prompt for the Visual Question Answering (VQA) model to help it identify the "{CRITERION}" category in a given image at three different granularity. Please help me design and generate this prompt using the following template: ”Question: [Generated VQA Prompt Question] Answer (reply with an abstract, a common, and a specific category name, respectively):”. The generated prompt should be simple and straightforward.
@@
In the tag-based grouper, we begin by embedding the target criterion into the prompt from Tab. 11 to generate criterion-specific tags at a middle granularity. To enhance the diversity and coverage of the tags, we query the LLM 10 times and take the union of the generated tags after deduplication as candidates. Following the SHiNe framework [liu2024shine], for each middle-grained tag, we further embed it into the prompts from Tab. 12 and Tab. 13 to generate 3 super-categories (coarse-grained) and 10 sub-categories (fine-grained) for each tag.
After generating coarse and fine-grained categories for all middle-grained tags, we take the union of the super-categories as the coarse-grained tag candidates and the union of the sub-categories as the fine-grained tag candidates. Lastly, we use the open-vocabulary tagger CLIP ViT-L/14 to assign the most relevant tags to each image based on cosine similarity, using candidates from each granularity level. After tagging all the images, we group those sharing the same tag into clusters, yielding the clustering result. Note that we do not utilize lexical databases such as WordNet [miller1995wordnet] or ConceptNet [speer2017conceptnet] for tag generation, as they do not support free-form input and may not capture certain discovered criteria.
@@
Output Instruction Please respond in the following format within five words: ”*Answer*”. Do not talk about the description and do not respond long sentences. The answer should be within five words.
@@
Sub-task Instruction Follow these steps to generate the hierarchy: Step 1 - Understand the provided initial list of "{CRITERION}" categories. The following three-level class hierarchy generation steps are all based on the provided initial list. Step 2 - Generate a list of abstract or general "{CRITERION}" categories as the first level of the class hierarchy, covering all the concepts present in the initial list. Step 3 - Generate a list of middle-grained "{CRITERION}" categories as the second level of the class hierarchy, in which the middle-grained categories are the subcategories of the categories in the first level. The categories in the second-level are more specific than the first level but should still cover and reflect all the concepts present in the initial list. Step 4 - Generate a list of more specific fine-grained "{CRITERION}" categories as the third level of the class hierarchy, in which the categories should reflect more specific "{CRITERION}" concepts that you can infer from the initial list. The categories in the third-level are subcategories of the second-level. Step 5 - Output the generated three-level class hierarchy as a JSON object where the keys are the level numbers and the values are a flat list of generated categories at each level, structured like: { “level 1”: [“categories”], “level 2”: [“categories”], “level 3”: [“categories”] }
@@
We first present the MLLM prompt used for LLaVA-NeXT-7B [liu2024llava] to generate criterion-specific captions in Tab. 14. Following this, we present the LLM Llama-3.1-8B prompts used in the caption-based grouper for the Initial Naming step in Tab. 15, the Multi-granularity Cluster Refinement step in Tab. 16, and the Final Assignment step in Tab. 17.
@@
Appendix G Further Details of the Compared Methods
@@
Implementation details of MMaP and MSub [yao2024multi, yao2025customized]: We closely followed the training configuration outlined in the original MMaP and MSub paper. Specifically, GPT-turbo-2024-04-09 was used as the LLM to generate reference words for each dataset. We then prompt-tuned CLIP-ViT/B32 using Adam with a momentum of 0.9, training the model for 1,000 epochs for each criterion across all datasets. Hyperparameters were optimized according to the loss score of MMaP, with the learning rate searched in {0.1, 0.05, 0.01, 0.005, 0.001, 0.0005}, weight decay in {0.0005, 0.0001, 0.00005, 0.00001, 0}, $\alpha$ and $\beta$ in {0.0, 0.1, 0.2, …, 1.0}, and $\lambda$ fixed at 1 for all experiments. After training, KMeans, with the ground-truth number of clusters, was applied for each criterion and dataset to perform clustering.
Appendix H Further Quantitative Experimental Results
In this section, we present additional numerical experiment results to supplement the figures in the main paper. In § H.1, we provide supplementary results for the evaluation of the Criteria Proposer in our framework. In § H.2, we present additional results for the evaluation of the Semantic Grouper across various criteria on the six tested benchmarks. Furthermore, we include expanded results comparing our framework to prior criteria-conditioned clustering methods. Lastly, we present detailed results from the ablation study of the multi-granularity refinement component in § H.4.
H.1 Further Results for Criteria Proposer Study
@@
Although captions generated by the MLLM may exhibit some information loss (e.g., ignoring small objects or attributes) [he2024incorporating] and hallucinations (e.g., introducing objects not present in the images) [liu2024survey], these issues generally occur at the object or fine-grained attribute level. However, when reasoning about grouping criteria for OpenSMC task, the focus is on identifying general thematic elements shared across the image set. As a result, these minor inconsistencies in the captions do not hinder the LLM in our framework from effectively reasoning about grouping criteria, helping the Caption-based Proposer to achieve the best performance among all the studied design choices.
@@
H.2 Further Results for Semantic Grouper Study
@@
H.3 Further Comparative Results with TCMC Methods
@@
H.4 Further Results for Multi-granularity Clustering Study.
@@
Appendix I Qualitative Analysis
@@
Appendix J Failure Case Analysis
@@
Finally, we observed that our method sometimes fails to distinguish subtle, fine-grained differences between images, leading to incorrect labels. For example, as shown in Figure 8, “Edamame” or “Pho” are typical dishes from China, Vietnam, and Japan, but they may be presented differently depending on the cuisine. The “Edamame” shown in Figure 8 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine. This oversight of fine-grained details could be improved by employing a more advanced prompting strategy [liu2024democratizing].
@@
Appendix K Further Study on Multi-granularity Clustering
@@
Appendix L Study on Handling Invalid Criteria
@@
To this end, we design and conduct a control experiment using the Fruit-2c dataset [muresan2018fruit], where we artificially introduced two “hallucinated” invalid grouping criteria (False Positives), Action and Clothing Style, into the refined criteria pool. These invalid criteria were then used in the subsequent grouping process to evaluate their effect on our system. We apply the main Caption-based Grouper to group fruit images based on these “hallucinated” criteria.
@@
Appendix M Study on Model Hallucination and Bias
@@
Model hallucination. LLM hallucination [wang2024factuality] typically occurs when LLMs are tasked with complex queries requiring world knowledge or factual information—for instance, answering a question like ”Who was the 70th president of the United States?” might lead to a fabricated response. However, in our system, the use of LLMs is fully grounded in the visual descriptions (tags or captions) of the images. Consequently, the LLM output is strongly constrained to analyzing these visual descriptions, significantly reducing the likelihood of hallucination. That said, LLM hallucination can still have mild effects on clustering results. For example, as discussed in the failure case analysis in Appendix J, the LLM incorrectly grouped “Korean bibimbap” and “Vietnamese rice noodles” under “Chinese cuisine” (see Figure 8). MLLMs also play a crucial role in our system, as they are responsible for translating images into text for subsequent processing steps. MLLM hallucination [wang2024factuality] typically involves incorrectly identifying the existence of objects, attributes, or spatial relationships within an image. However, since our proposed system operates at the dataset level rather than on a per-image basis, it is largely insensitive to such hallucinations, especially at the fine-grained visual detail level. Moreover, as our system is training-free, it can be further enhanced with LLM or MLLM hallucination mitigation techniques, such as the Visual Fact Checker [ge2024visual], which we leave as a direction for future work.
@@
Model Bias. Foundation models such as LLMs and MLLMs are known to inherit biases from their training data [bommasani2021opportunities]. In our system, we addressed potential biases using Hard Positive Prompting techniques: i) MLLM Bias Mitigation: The MLLM is further prompted to generate criterion-specific captions that focus solely on describing the criterion-related content in each image. This approach constrains the MLLM from generating irrelevant content influenced by inherent biases; ii) LLM Bias Mitigation: Similarly, when prompting the LLM to assign image captions to clusters, we condition it to concentrate exclusively on the Criterion depicted in each image (see Tab. 15).
@@
To validate the effectiveness of these bias mitigation techniques, we conducted a fair clustering experiment. Specifically, following kwon2023image, we sampled images for four occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET [gustafson2023facet] dataset, which contains images from 52 occupations. For each occupation, we selected 10 images of men and 10 images of women, totaling 80 images, ensuring a ground-truth gender proportion disparity of 0% for each occupation. Using our main $\mathcal{X}$-Cluster system, we grouped these images based on the criterion Occupation using three bias mitigation strategies: i) No mitigation: using general descriptions from the MLLM for LLM grouping; ii) Our default hard positive prompting strategy: using criterion-specific captions from the MLLM for LLM grouping; and iii) Our default strategy with additional negative prompt: adding a simple negative prompt, “Do not consider gender,” to both the MLLM captioning and LLM grouping prompts.
@@
Appendix N Computational Cost Analysis
The proposed $\mathcal{X}$-Cluster framework is training-free, requiring only inference processes. Specifically, our main framework (Caption-based) requires up to 31 GB of GPU memory to operate. All experiments reported in the paper were conducted on 4 Nvidia A100 40GB GPUs. In Tab. 31, we provide a detailed analysis of the computational efficiency of our main $\mathcal{X}$-Cluster framework (Caption-based Proposer and Caption-based Grouper) on the COCO-4c benchmark (5,000 images with four criteria) across various hardware configurations. For these experiments, we used LLaVA-NeXT-7B [liu2024llava] as the MLLM and Llama-3.1-8B [meta2024llama31] as the LLM.
[…]
@@
Appendix O System Sensitivity Analysis of Various MLLMs and LLMs
@@
Specifically, in Figure 10(a), we first fix the LLM in our system to Llama-3.1-8B [meta2024llama31] and assess the influence of various MLLMs: GPT-4V [openai2023gpt4], BLIP-3-4B [xue2024xgen], and LLaVA-NeXT-7B [liu2024llava]. Next, in Figure 10(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo [openai2023gpt4], GPT-4o [openai2024gpt4o], Llama-3-8B [meta2024llama3], and Llama-3.1-8B.
@@
As for the LLMs, as depicted in Figure 10(b), despite GPT-4-turbo showing marginally superior performance, the open-source Llama-3.1-8B achieves similar results across benchmarks, making it our default LLM. Notably, except for the Card-2c dataset, system performance remains largely consistent regardless of the power of the LLM. This consistency suggests that the reasoning task for OpenSMC, given the capabilities of modern LLMs to tackle complex problems [street2024llms], is relatively straightforward.
Appendix P Study on Fine-grained Image Collections
Image collections may include fine-grained grouping criteria, such as Bird species in bird photography. Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations [zhang2014part, vedaldi2014understanding, he2017fine]. This requires the model to detect subtle visual distinctions to accurately infer cluster names and guide the grouping process. The straightforward captioning process in our current framework may not fully capture these subtle visual nuances. However, the modular design of our framework allows for seamless integration of advanced cross-modal chain-of-thought (CoT) prompting strategies to address this issue.
We demonstrate this by enhancing our Caption-based Grouper with FineR [liu2024democratizing], a cross-modal CoT prompt method specifically designed for fine-grained visual recognition. When the proposer identifies fine-grained criteria, such as Bird species, the framework switches to a FineR-enhanced captioning strategy that provides more detailed attribute descriptions, such as “Wing color: Blue-grey,” to enrich the captions and capture per-attribute visual characteristics to better support the subsequent substructure uncovering process.
Table 32: Study of substructure discovery for fine-grained criteria. We report clustering accuracy (CAcc) and semantic accuracy (SAcc) as percentages (%). The pseudo upper-bound (UB) performance is obtained using CLIP [radford2021learning] ViT-L/14 in a zero-shot transfer setting with the ground-truth class names. †: We compare with FineR [liu2024democratizing] without its post-class name refinement step to ensure a fair comparison.